# FP4 All the Way: Fully Quantized Training of LLMs

### v1

@@ 6 Discussion
This work presents the first demonstration of fully quantized FP4 training—covering weights, activations, and gradients—at large scale. Our experiments on Llama2 7B model show that, while a small gap in training loss initially appears compared to BF16, this gap can be fully closed with a short QAF phase where the forward pass remains in FP4 and only the backward pass switches to BF16. Importantly, downstream task performance remains on par with BF16, confirming FP4’s practical viability.
@@
A key limitation of this work is the lack of dedicated FP4 support in current Gaudi hardware, which prevents us from directly measuring the potential speedup and energy efficiency benefits of native FP4 execution. As a result, all experiments are conducted using FP4 simulations in Gaudi2, which incur additional overhead from precision casting and lead to longer runtimes. Based on previous FP8 works [13, 9] we expect in a rough estimation to $\sim 35-40\%$ time-to-train acceleration in comparison to FP8, which corresponds to $\sim 85\%$ time-to-train acceleration in comparison to the BF16 baseline.
@@ B.1 Analysis of quantized SGD with stochastic rounding
$-\eta\,\nabla L(\theta_{t})^{T}\nabla L(\theta_{t})=-\eta\,\|\nabla L(\theta_{t})\|^{2}.$
@@
$\mathbb{E}[L(\theta_{t+1})-L(\theta_{t})]\approx-\eta\|\nabla L(\theta_{t})\|_{2}^{2}+\tfrac{1}{2}\eta^{2}\left(\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})+\sigma_{q}^{2}\mathrm{tr}(H(\theta_{t}))\right)$
@@
$\mathbb{E}[L(\theta_{t+1})-L(\theta_{t})]\approx\underbrace{-\left(\eta\|\nabla L(\theta_{t})\|_{2}^{2}-\tfrac{1}{2}\eta^{2}\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})\right)}_{\text{useful descent component}}+\underbrace{\tfrac{1}{2}\eta^{2}\sigma_{q}^{2}\mathrm{tr}(H(\theta_{t}))}_{\text{quantization noise effect}}$
@@
$\eta\|\nabla L(\theta_{t})\|_{2}^{2}>\tfrac{1}{2}\eta^{2}\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})\Rightarrow\eta<\frac{2\|\nabla L(\theta_{t})\|_{2}^{2}}{\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})}$
@@
$\eta<\frac{2}{\lambda_{\max}(H(\theta_{t}))}$
@@
$U(\eta)=\mathbb{E}[L(\theta_{t+1})-L(\theta_{t})]=-\eta\|\nabla L(\theta_{t})\|_{2}^{2}+\tfrac{1}{2}\eta^{2}\left(\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})+\sigma_{q}^{2}\mathrm{tr}(H(\theta_{t}))\right)$
Set derivative to 0:
$\frac{dU}{d\eta}=-\|\nabla L(\theta_{t})\|_{2}^{2}+\eta\left(\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})+\sigma_{q}^{2}\mathrm{tr}(H(\theta_{t}))\right)=0$
@@
$\eta^{*}=\frac{\|\nabla L(\theta_{t})\|_{2}^{2}}{\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})+\sigma_{q}^{2}\mathrm{tr}(H(\theta_{t}))}$
@@
$\frac{df}{d\sigma_{q}}=Z\,Y\;\frac{(X+Y\sigma_{q}^{2})^{2}\;-\;\sigma_{q}\cdot 2\,(X+Y\sigma_{q}^{2})\cdot(2Y\sigma_{q})}{(X+Y\sigma_{q}^{2})^{4}}$
Factor and simplify:
@@
$\sigma_{\mathrm{critical}}^{2}=\frac{\nabla L(\theta_{t})^{T}H(\theta_{t})\,\nabla L(\theta_{t})}{3\,\mathrm{tr}\!\bigl{(}H(\theta_{t})\bigr{)}}$
Finally, assuming [^1] In the high-dimensional regime ($d,N\to\infty$ and $d/N\rightarrow\lambda$, where $N$ is number of training samples), previous works [14] showed, using random-matrix theory and empirical results, that the bulk of the Hessian values can be approximately represented by the Marchenko–Pastur distribution, especially for small loss values. Thus, when $\lambda\ll 1$ (which is the common regime for LLMs), this Marchenko–Pastur distribution implies that the bulk of Hessian eigenvalues concentrates near their mean $\mathrm{tr}(H)/d$. Empirical investigations [16, 11] confirm that the gradient predominantly occupies this bulk subspace rather than aligning with the few extreme modes. Hence the curvature experienced in the gradient direction approximates the average eigenvalue.
@@
In other words, once the average per-coordinate gradient falls to $\sqrt{3}$ times the quantization-noise std, FP4 gradients lose efficacy and it is time to switch to higher precision. As shown in Figure 4, setting the noise std to $k\cdot\sigma_{\mathrm{critical}}$ with $k=2.0,1.0,0.5$ yields markedly different convergence behaviors around that threshold. In Section B.2 we show a similar analysis without SR, which shows that the “useful descent” vanishes to zero as we train, while the biased-noise term remains even after long training..

### v2

@@ 6 Discussion
This work presents the first demonstration of fully quantized FP4 training—covering weights, activations, and gradients—at large scale. Our experiments on Llama2 7B model show that, while a small gap in training loss initially appears compared to BF16, this gap can be fully closed with a short QAF phase, where the forward pass remains in FP4 and only the backward pass switches to BF16. Importantly, downstream task performance remains on par with BF16, confirming FP4’s practical viability.
@@
A key limitation of this work is the lack of dedicated FP4 support in current Gaudi hardware, which prevents us from directly measuring the potential speedup and energy efficiency benefits of native FP4 execution. As a result, all experiments are conducted using FP4 simulations in Gaudi2, which incur additional overhead from precision casting and lead to longer runtimes. Based on previous FP8 works (Peng et al., 2023; Fishman et al., 2024) we expect in a rough estimation to $\sim 35-40\%$ time-to-train acceleration in comparison to FP8, which corresponds to $\sim 85\%$ time-to-train acceleration in comparison to the BF16 baseline. Our work centers on the LLaMA architecture, one of the most widely adopted frameworks in modern LLMs. Preliminary experiments indicate that the approach can be directly applied to Mixture-of-Experts (MoE) architectures. Further analysis of MoE models and extensions to vision tasks are reserved for future work.
@@ B.1 Analysis of quantized SGD with stochastic rounding
$\eta\,\nabla L(\theta_{t})^{T}\nabla L(\theta_{t})=-\eta\,\|\nabla L(\theta_{t})\|^{2}.$
@@
$\mathbb{E}[L(\theta_{t+1})-L(\theta_{t})]\approx-\eta\|\nabla L(\theta_{t})\|_{2}^{2}+\tfrac{1}{2}\eta^{2}\left(\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})+\sigma_{q}^{2}\mathrm{tr}(H(\theta_{t}))\right)\,.$
@@
$\mathbb{E}[L(\theta_{t+1})-L(\theta_{t})]\approx\underbrace{-\left(\eta\|\nabla L(\theta_{t})\|_{2}^{2}-\tfrac{1}{2}\eta^{2}\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})\right)}_{\text{useful descent component}}+\underbrace{\tfrac{1}{2}\eta^{2}\sigma_{q}^{2}\mathrm{tr}(H(\theta_{t}))}_{\text{quantization noise effect}}\,.$
@@
$\eta\|\nabla L(\theta_{t})\|_{2}^{2}>\tfrac{1}{2}\eta^{2}\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})\Rightarrow\eta<\frac{2\|\nabla L(\theta_{t})\|_{2}^{2}}{\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})}\,.$
@@
$\eta<\frac{2}{\lambda_{\max}(H(\theta_{t}))}\,.$
@@
To find the optimal step size, we define
$U(\eta)=\mathbb{E}[L(\theta_{t+1})-L(\theta_{t})]=-\eta\|\nabla L(\theta_{t})\|_{2}^{2}+\tfrac{1}{2}\eta^{2}\left(\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})+\sigma_{q}^{2}\mathrm{tr}(H(\theta_{t}))\right)\,.$
Setting its derivative to 0:
$\frac{dU}{d\eta}=-\|\nabla L(\theta_{t})\|_{2}^{2}+\eta\left(\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})+\sigma_{q}^{2}\mathrm{tr}(H(\theta_{t}))\right)=0\,.$
@@
$\eta^{*}=\frac{\|\nabla L(\theta_{t})\|_{2}^{2}}{\nabla L(\theta_{t})^{T}H(\theta_{t})\nabla L(\theta_{t})+\sigma_{q}^{2}\mathrm{tr}(H(\theta_{t}))}\,.$
@@
$\frac{df}{d\sigma_{q}}=Z\,Y\;\frac{(X+Y\sigma_{q}^{2})^{2}\;-\;\sigma_{q}\cdot 2\,(X+Y\sigma_{q}^{2})\cdot(2Y\sigma_{q})}{(X+Y\sigma_{q}^{2})^{4}}\,.$
We factor this expression and simplify it:
@@
$\sigma_{\mathrm{critical}}^{2}=\frac{\nabla L(\theta_{t})^{T}H(\theta_{t})\,\nabla L(\theta_{t})}{3\,\mathrm{tr}\!\bigl{(}H(\theta_{t})\bigr{)}}\,.$
Finally, assuming [^1] In the high-dimensional regime ($d,N\to\infty$ and $d/N\rightarrow\lambda$, where $N$ is number of training samples), previous works Pennington and Worah [2017] showed, using random-matrix theory and empirical results, that the bulk of the Hessian values can be approximately represented by the Marchenko–Pastur distribution, especially for small loss values. Thus, when $\lambda\ll 1$ (which is the common regime for LLMs), this Marchenko–Pastur distribution implies that the bulk of Hessian eigenvalues concentrates near their mean $\mathrm{tr}(H)/d$. Empirical investigations Sagun et al. [2017], Ghorbani et al. [2019] confirm that the gradient predominantly occupies this bulk subspace rather than aligning with the few extreme modes. Hence the curvature experienced in the gradient direction approximates the average eigenvalue.
@@
In other words, once the average per-coordinate gradient falls to $\sqrt{3}$ times the quantization-noise std, FP4 gradients lose efficacy and it is time to switch to higher precision. As shown in Figure 4, setting the noise std to $k\cdot\sigma_{\mathrm{critical}}$ with $k=2.0,1.0,0.5$ yields markedly different convergence behaviors around that threshold. In Section˜B.2 we show a similar analysis without SR, which shows that the “useful descent” vanishes to zero as we train, while the biased-noise term remains even after long training..
@@ B.2 Impact of nonzero mean noise (deterministic rounding)
Appendix C Additional Experimental Results
Figure 7: Comparison of different rounding schemes when training a 350M Llama model using NVFP4 format. In each graph, we apply RtN in one of the six elements in one of the GEMMs while the rest use SR. Notice that applying RtN to neural gradients during both the ‘Update’ and ‘Backward’ GEMMs, and to the activations during the ‘Update’ GEMM leads to higher training loss, while applying RtN to the other components has the opposite effect, reducing the loss.
Table 4: Training loss of Llama 125M over 30B tokens with different seeds, yield a similar loss, resulting in a standard deviation of 0.001.
Training loss Seed
3.03 1337
3.027 1234
3.025 2345
3.027 3456
3.029 4567
Table 5: Ablation study to determine how many tokens are required to reach the same loss as the BF16 baseline. Notice the QAF ratio decrease for larger dataset. In all QAF experiments we use the peak learning rate equal to the last learning rate in the FP4 training.
Starting point QAF length Ratio
200B 20B 10%
500B 28B 5.6%
1T 40B 4%